OpenStudy (anonymous):

Mardi received an inheritance of $60,000. She invested part at 10% and deposited the rest the remainder in tax-free bonds at 11%. Her total annual income from the investment was $6100. Fine the amount invested at 10%

OpenStudy (anonymous):

A*10% + B*11%=6100 Where A is the investment and B is the deposit

OpenStudy (anonymous):

We know that A + B= 60,000

OpenStudy (anonymous):

We want to know what A is right ? So we want to replace the B in the first equation with an A. From the second equation I gave you A + B= 60,000 B= 60,000-A

OpenStudy (anonymous):

Replace that with the first and solve for A

OpenStudy (anonymous):

A*0.10 + B*0.11=6100

OpenStudy (anonymous):

thank you. answer is 50,000 :)